The total amount will be rs 180.If each of student contribute rs x and our teacher contribute rs 100 .express the statement in the form of linear quadratic equation with one variable.
Answer by Theo(13342)

180 = 100 + y * x

y is the number of students.
x is the amount that each student contributes.

for example, if there are 10 students, then the formula becomes 180 = 100 + 10 * x

you would solve for x to get x = 8.

the teacher contributes 100, and 10 students contribute 8 apiece for a total of 180.

if you don't know the number of students, nor what each student needs to contribute, then your formula remains:

180 = 100 + y * x.

if you know what each student needs to contribute, then you can solve for y which tells you the number of students you need.

if you know the number of students, then you can solve for x which tells you what each student needs to contribute.

you cannot have just one variable unless you either know the number of students or you know the amount that each student needs to contribute.